OpenStudy (anonymous):

If x(5x -4)(x+3)(7x-2)= Ax^4 +Bx^3 +Cx^2 +Dx+E, find A+B+C+D+E. Please show me all of your work.

myininaya (myininaya):

you should multiply all of that scary stuff out \[x(5x^2+15x-4x-12)(7x-2)=x(5x^2+11x-12)(7x-2)=\] \[=x(35x^3-10x^2+77x^2-22x-84x+24)=x(35x^3+66x^2-106x+24)\] \[=35x^4+66x^3-106x^2+24x=35x^4+66x^3-106x^2+24x+0\]

myininaya (myininaya):

\[35x^4+66x^3-106x^2+24x+0=Ax^4+Bx^3+Cx^2+Dx+E\]

myininaya (myininaya):

so what is A B C D E ?

OpenStudy (anonymous):

A=35 B=66 C=-106 D=24 E=0 Is it right

myininaya (myininaya):

and then add them to find that sum! :)

OpenStudy (anonymous):

Is it 19

myininaya (myininaya):

35+66=101 + 24 ------------ 125 and 125-106=19 yep thats what I get :)

OpenStudy (anonymous):

I need to do a lot of practice. Thanks a lot for helping me.

myininaya (myininaya):

no problem. i showed all the steps i would take in order to solve this problem. i hope it was helpful.

OpenStudy (anonymous):

yes.It is.

myininaya (myininaya):

:) have a good day!

OpenStudy (anonymous):

Thanks again